# DRIFT-Net: A Spectral–Coupled Neural Operator for PDEs Learning

### v1

@@ Abstract
Learning PDE dynamics with neural solvers can significantly improve wall-clock efficiency and accuracy compared with classical numerical solvers. In recent years, foundation models for PDEs have largely adopted multi-scale windowed self-attention, with the scOT backbone in Poseidon serving as a representative example. However, because of their locality, truly globally consistent spectral coupling can only be propagated gradually through deep stacking and window shifting. This weakens global coupling and leads to error accumulation and drift during closed-loop rollouts. To address this, we propose DRIFT-Net. It employs a dual-branch design comprising a spectral branch and an image branch. The spectral branch is responsible for capturing global, large-scale low-frequency information, whereas the image branch focuses on local details and nonstationary structures. Specifically, we first perform controlled, lightweight mixing within the low-frequency range. Then we fuse the spectral and image paths at each layer via bandwise weighting, which avoids the width inflation and training instability caused by naive concatenation. The fused result is transformed back into the spatial domain and added to the image branch, thereby preserving both global structure and high-frequency details across scales. Compared with strong attention-based baselines, DRIFT-Net achieves lower error and higher throughput with fewer parameters under identical training settings and budget. On Navier–Stokes benchmarks, the relative $L_{1}$ error is reduced by 7%–54%, the parameter count decreases by about 15%, and the throughput remains higher than scOT. Ablation studies and theoretical analyses further demonstrate the stability and effectiveness of this design. The code is available at https://github.com/cruiseresearchgroup/DRIFT-Net.
@@ 2 Related Work
Neural operators learn PDE solution mappings directly between function spaces. The Fourier Neural Operator (FNO) introduced spectral-domain convolutions to efficiently capture global interactions  (Li et al., 2021), while DeepONet adopts a branch and trunk factorization grounded in operator approximation theory  (Lu et al., 2021). Building on these paradigms, several variants have broadened the applicability of neural operators. AFNO leverages low-rank factorization for high-resolution efficiency  (Guibas et al., 2022). U-NO deepens the U-Net style encoder–decoder hierarchy for multi-scale representations  (Rahman et al., 2022). Geo-FNO extends spectral operators to irregular geometries  (Li et al., 2023), and CNO targets robust multi-scale approximation in challenging settings  (Raonić et al., 2023). Beyond single-task settings, recent work explores foundation models for scientific ML to improve scaling and transfer (Subramanian et al., 2023).
Foundation models with multi-scale attention.
Within PDE foundation models, Poseidon employs a multi-scale windowed self-attention operator Transformer (scOT) with a U-Net hierarchy and time-conditioned normalization. It demonstrates strong transfer performance across heterogeneous equation families  (Herde et al., 2024; Liu et al., 2021; 2022). Windowed self-attention partitions the domain into local regions to compute attention efficiently, but its locality implies that domain-wide dependencies are established only gradually via deep stacking and window shifts. Empirically, this can weaken globally consistent spectral coupling and induce rollout drift over long horizons  (Lippe et al., 2023). In practice, naive cross-scale or cross-branch concatenation may inflate channel width and destabilize training, whereas purely spectral operators, while global, often underfit nonstationary local details.
@@
DRIFT-Net augments a multi-scale attention backbone with an explicit spectral path for controlled low-frequency mixing. It also introduces a non-expansive, bandwise fusion mechanism that integrates the spectral and image branches without inflating the feature width. Methodological details and quantitative comparisons appear in Sections  4 and  5.
@@ 4.2 Controlled Low-Frequency Mixing
For later use we denote $\hat{V}_{\text{low}}(k)\!=\!\mathbb{1}_{k\in M_{\text{low}}}\hat{V}(k)$ and keep $\hat{X}_{\text{high}}(k)\!=\!\hat{X}(k)-\mathbb{1}_{k\in M_{\text{low}}}\hat{X}(k)$. In practice, $W$ is unconstrained (no explicit spectral-norm clipping), which keeps the transform expressive. We use rFFT2/iFFT2 and the inverse transform yields a real-valued field by construction (Cooley & Tukey, 1965; Frigo & Johnson, 2005). This design focuses global mixing on large scales while leaving high-frequency content intact. For implementation details, refer to Appendix C.
@@ 5 Experiments
We follow the POSEIDON evaluation protocol Herde et al. (2024) and compare DRIFT-Net against two strong baselines: SCOT (multi-scale windowed self-attention) and FNO (a Fourier neural operator). All models use the same training/validation/test split and preprocessing. They are trained with an identical data budget (approximately 20,000 trajectories over 40 epochs) and identical optimization settings (the same optimizer, learning-rate schedule, batch size, numerical precision, and hardware). On a discrete grid with $C$ channels and $H\times W$ spatial points, we train a one-step operator $F_{\theta}:\,u_{t}\mapsto u_{t+1}$ using teacher forcing. At test time, we apply $F_{\theta}$ autoregressively in a closed loop to reach the common target time $T^{*}$. We report the test-set mean relative $L^{1}$ error at the final time:
@@
We evaluate DRIFT-Net on four canonical unsteady Navier–Stokes benchmarks from the POSEIDON suite Herde et al. (2024). NS-SL (shear layer) tests vortex roll-up and mixing from a perturbed interface. NS-PwC (piecewise-constant vorticity) stresses the advection of sharp discontinuities. NS-Tracer-PwC extends NS-PwC by introducing a passive scalar $c$, which requires accurate coupling between the velocity and tracer fields. FNS-KF (forced Kolmogorov flow) sustains two-dimensional turbulence via steady forcing, challenging long-horizon stability and multi-scale fidelity.
In addition to these four benchmarks, we also consider two ApeBench-generated forced Kolmogorov variants constructed using a high-accuracy pseudo-spectral solver Koehler et al. (2024). These two variants share the same PDE setup and preprocessing but differ in physical parameters: one is more turbulent, whereas the other is comparatively smoother. To emphasize long-horizon robustness, we evaluate on both variants with a closed-loop rollout length of $T=100$ and, for these two datasets, compare DRIFT-Net against SCOT under identical settings.
@@
Table 1 reports the final-time relative $L^{1}$ errors under matched training conditions and evaluation protocol. On all four POSEIDON benchmarks, DRIFT-Net attains the lowest error among the compared methods. The table also includes the two long-horizon ApeBench-based Kolmogorov variants (rollout to T=100T{=}100): on the turbulent variant, DRIFT-Net achieves a lower final error than SCOT; on the smoother variant, DRIFT-Net again yields a lower final error, indicating improved robustness across physical regimes. For completeness, we further summarize the mean error across the rollout and the linear growth rate in the paragraph following the table, and present the corresponding error–time curves in Figure 2.
Table 1: Final-time relative $L^{1}$ error (lower is better). Closed-loop rollouts to the common target time $T^{*}$ for POSEIDON benchmarks, and to T=100T{=}100 for our two ApeBench-based long-horizon Kolmogorov datasets. Values are test-set means. Best in bold.
Task SCOT FNO DRIFT-Net
NS-SL (POSEIDON) 3.96 3.69 3.40
NS-PwC (POSEIDON) 2.35 4.57 1.09
NS-Tracer-PwC (POSEIDON) 5.18 9.46 4.19
FNS-KF (POSEIDON) 4.65 4.43 4.32
@@
For the two ApeBench-based datasets, we also report the mean relative $L^{1}$ error across the rollout and a linear growth slope obtained by a least-squares fit of error versus time. On the turbulent variant, DRIFT-Net attains a mean error of $69.89$ compared with $74.76$ for SCOT, and an error growth slope of $1.154$ compared with $1.185$ for SCOT. On the smoother variant, the mean error is $31.07$ compared with $35.41$ for SCOT, and the slope is $0.581$ compared with $0.641$ for SCOT. Figure 2 depicts the corresponding error–time curves (left: turbulent; right: smoother), showing consistently lower curves and flatter tails for DRIFT-Net.
Long-horizon behavior.
Although final-time metrics are informative, the full closed-loop trajectory over a long horizon provides additional insight into cumulative drift. Figure 3 presents the test-set mean relative $L^{1}$ error at each time step for NS-Tracer-PwC. DRIFT-Net achieves a final-time error of $4.193$, whereas SCOT achieves $5.182$ (a reduction of $19.1\%$). The average error across the rollout is $1.99$ for DRIFT-Net and $2.46$ for SCOT, and the linear growth slopes are $0.176$ and $0.229$, respectively. These results are consistent with the objective of strengthening global low-$k$ coupling while maintaining high-$k$ fidelity.
@@
Figure 3: Error vs. time on NS-Tracer-PwC. Solid line with circles: DRIFT-Net; dashed line with squares: SCOT. We plot the test-set mean relative $L^{1}$ at each time step under closed-loop rollouts.
Additional final-time results (subset) and context.
For the tasks where error time-series are available for both models, the final-time results confirm the advantage of DRIFT-Net over SCOT (Table 2). For completeness, we also include NS-PwC, on which DRIFT-Net achieves a final error of 1.090 compared with $2.350$ for SCOT (as reported in prior results).
Table 2: Additional final-time results (subset). Final-time relative $L^{1}$ (lower is better) where both series are available, and percent improvement of DRIFT-Net vs. SCOT. Best in bold.
Task DRIFT-Net ($e_{\text{rel-L1}}@T^{*}$) SCOT ($e_{\text{rel-L1}}@T^{*}$) % Improvement
NS-Tracer-PwC 4.193 5.182 19.119.1%
FNS-KF 4.315 4.647 7.27.2%
NS-SL 3.402 3.956 14.014.0%
NS-PwC 1.090 2.350 53.653.6%
@@
We also evaluate model size and inference speed under identical inference conditions (same hardware, numerical precision, batch size, and sequence length; excluding data-loading overhead). DRIFT-Net contains approximately 17 million parameters (about 15% fewer than SCOT’s 20 million) and achieves roughly 158 time-step predictions per second, whereas SCOT achieves about 118 steps/s (Table 3). The Conclusion reports 158 vs. 118 steps/s under a different resolution and precision configuration; both measurements reflect a similar relative advantage for DRIFT-Net, and we include both for completeness.
Table 3: Model size and inference throughput (higher throughput is better). Throughput measured under identical inference settings.
Model Parameters (M) Throughput (steps/s)
DRIFT-Net 17 158
SCOT 20 118
@@
We assess the contribution of each major component of DRIFT-Net: Low-Frequency Mixing (LFM), Radial Gating (RG), and Frequency-Weighted Loss (FWL). We train ablated variants on NS-PwC and NS-Tracer-PwC, holding all other hyperparameters fixed and using the same training budget and evaluation protocol as the full model. Removing LFM or RG noticeably worsens final-time accuracy; removing FWL also degrades performance, although to a smaller degree (Table 4). Specifically, relative to the full model, eliminating LFM increases the final error by $0.56$ on NS-PwC and by $2.13$ on NS-Tracer-PwC. Removing RG increases the error by $0.61$ and $2.36$ on these tasks, respectively. Dropping FWL yields smaller increases of $0.27$ and $1.17$.
@@ 6 Limitations and Future Work
DRIFT-Net has limitation. It uses a mask to select low frequencies. The initial band is hand tuned. The mask learns weights later, but the cutoff can be task specific. The model applies FFTs in many blocks. This adds memory traffic and latency on very large grids. Our tests target two dimensional flow. Three dimensional flow and coupled multi-physics may bring new training issues and higher cost. Future work will learn spectral partitions end to end. We will study 3D and multi-physics PDEs. We will pair the model with adaptive resolution and mesh refinement. We will also test irregular domains and complex boundary conditions.
@@ 7 Conclusion
We introduced DRIFT-Net for operator learning. The model couples a spectral branch with an image branch. It applies controlled low frequency mixing in the spectrum. It fuses branches with bandwise radial gating in a stable way. A frequency weighted loss reduces spectral bias. On Navier–Stokes benchmarks, DRIFT-Net lowers final time relative L1 error across all tasks. The reductions range from 7% to 54%. The model uses about 15% fewer parameters than SCOT and keeps higher throughput. For example, it reaches 158 steps per second while SCOT reaches 118 under the same setup. The unit is modular and can replace windowed attention blocks in existing backbones.

### v2

@@ Abstract
Learning PDE dynamics with neural solvers can significantly improve wall-clock efficiency and accuracy compared with classical numerical solvers. In recent years, foundation models for PDEs have largely adopted multi-scale windowed self-attention, with the scOT backbone in Poseidon serving as a representative example. However, because of their locality, truly globally consistent spectral coupling can only be propagated gradually through deep stacking and window shifting. This locality can weaken globally consistent spectral coupling, which has been associated with increased error accumulation and drift during closed-loop rollouts. To address this, we propose DRIFT-Net. It employs a dual-branch design comprising a spectral branch and an image branch. The spectral branch is responsible for capturing global, large-scale low-frequency information, whereas the image branch focuses on local details and nonstationary structures. Specifically, we first perform controlled, lightweight mixing within the low-frequency range. Then we fuse the spectral and image paths at each layer via bandwise weighting, which avoids the width inflation and training instability caused by naive concatenation. The fused result is transformed back into the spatial domain and added to the image branch, thereby preserving both global structure and high-frequency details across scales. Compared with strong attention-based baselines, DRIFT-Net achieves lower error and higher throughput with fewer parameters under identical training settings and budget. On Navier–Stokes benchmarks, the relative $L_{1}$ error is reduced by 7%–54%, the parameter count decreases by about 15%, and the throughput remains higher than scOT. Ablation studies and theoretical analyses further demonstrate the stability and effectiveness of this design. The code is available at https://github.com/cruiseresearchgroup/DRIFT-Net.
@@ 2 Related Work
Neural operators learn mappings between function spaces and approximate solution operators of PDEs (Li et al., 2021; Kovachki et al., 2023). A recent survey categorizes existing architectures into three main types: deep operator networks (DeepONets), integral-kernel neural operators, and transformer-based neural operators (Liu et al., 2025). DeepONet adopts a branch–trunk factorization and comes with operator-approximation guarantees (Lu et al., 2021). Integral-kernel neural operators model nonlocal interactions through parameterized kernels. FNO implements the kernel operator in the Fourier domain (Li et al., 2021), and variants such as Geo-FNO and U-FNO extend this idea to irregular geometries and multiphase flow (Li et al., 2023; Wen et al., 2022); CNO and neural operators with localized integral or differential kernels further emphasize multi-scale structure and local operators (Raonić et al., 2023; Liu-Schiaffini et al., 2024). Transformer-based neural operators treat grid points as tokens and use self-attention to approximate data-dependent kernels, and are widely used in multi-task and pretraining settings (McCabe et al., 2023; Hao et al., 2024; Yang et al., 2023; Rahman et al., 2024).
PDE foundation models and multi-scale attention.
A recent line of work builds PDE foundation models on top of transformer-based neural-operator backbones. Examples include MPP (a spatio-temporal transformer surrogate trained autoregressively on multiphysics datasets), the operator transformer DPOT with Fourier attention, and attention-based neural operators such as ICON and CoDA-NO; all of them perform large-scale pretraining on multi-dataset or multiphysics corpora and transfer to downstream PDE tasks (McCabe et al., 2023; Hao et al., 2024; Yang et al., 2023; Rahman et al., 2024). Within multi-scale attention frameworks, Poseidon employs a multi-scale windowed attention operator transformer (scOT) with a U-Net hierarchy and time-conditioned normalization, and shows strong cross-task generalization on Navier–Stokes benchmarks (Herde et al., 2024; Liu et al., 2021; 2022). Windowed self-attention computes attention efficiently in local windows, but global dependencies are established only gradually through depth and window shifts, which can weaken global spectral coupling and contribute to rollout drift at long horizons (Lippe et al., 2023).
@@
We propose DRIFT-Net as a spectral–spatial coupled integral-kernel neural-operator backbone, which in the above taxonomy falls into the class of integral-kernel neural operators. Architecturally, DRIFT-Net is composed of stacked DRIFT blocks: each block augments the multi-scale attention backbone with an explicit spectral path for controlled low-frequency mixing and a non-expansive bandwise fusion mechanism that integrates the spectral and image branches without increasing the feature width. Methodological details and quantitative comparisons are given in Sections 4 and 5.
@@ 4.2 Controlled Low-Frequency Mixing
For later use we denote $\hat{V}_{\text{low}}(k)\!=\!\mathbb{1}_{k\in M_{\text{low}}}\hat{V}(k)$ and keep $\hat{X}_{\text{high}}(k)\!=\!\hat{X}(k)-\mathbb{1}_{k\in M_{\text{low}}}\hat{X}(k)$. In practice, $W$ is unconstrained (no explicit spectral-norm clipping), which keeps the transform expressive. We use rFFT2/iFFT2 and the inverse transform yields a real-valued field by construction (Cooley and Tukey, 1965; Frigo and Johnson, 2005). This design focuses global mixing on large scales while leaving high-frequency content intact. For implementation details, refer to Appendix C.
@@ 5 Experiments
We follow the POSEIDON evaluation protocol (Herde et al., 2024). All models, including DRIFT-Net and all baselines, use the same train/validation/test splits and preprocessing, and are trained under a shared data budget and training protocol: the same number of supervised trajectories per task, identical epoch budgets, and a common optimizer and learning-rate schedule inherited from the POSEIDON scOT configuration (see Appendix F.4). On a discrete grid with $C$ channels and $H\times W$ spatial points, we train a one-step operator $F_{\theta}:\,u_{t}\mapsto u_{t+1}$ with single-step teacher forcing. At test time, we apply $F_{\theta}$ autoregressively in a closed loop to reach the common target time $T^{*}$. We report the test-set mean relative $L^{1}$ error at the final time:
@@
Baselines.
We compare DRIFT-Net with scOT (Herde et al., 2024), FNO (Li et al., 2021), U-NO (Rahman et al., 2022), a U-Net–ConvNeXt baseline, and a Refiner U-Net baseline adapted from PDE-Refiner (Lippe et al., 2023); see Appendix for detailed configurations.
@@
We evaluate DRIFT-Net on four canonical unsteady Navier–Stokes benchmarks from the POSEIDON suite (Herde et al., 2024). NS-SL (shear layer) tests vortex roll-up and mixing from a perturbed interface. NS-PwC (piecewise-constant vorticity) stresses the advection of sharp discontinuities. NS-Tracer-PwC extends NS-PwC by introducing a passive scalar $c$, which requires accurate coupling between the velocity and tracer fields. FNS-KF (forced Kolmogorov flow) sustains two-dimensional turbulence via steady forcing, challenging long-horizon stability and multi-scale fidelity.
To cover different types of PDEs, we additionally include three tasks: Poisson–Gauss describes a stationary Poisson equation with Gaussian forcing and assesses performance on elliptic problems. The Allen-Cahn equation (ACE) models interface evolution and phase separation, representing a prototypical reaction-diffusion process. Wave-Gauss simulates two-dimensional wave propagation in a Gaussian medium and tests the modeling of hyperbolic wave phenomena.
In addition to these four benchmarks, we also consider two ApeBench-generated forced Kolmogorov variants constructed using a high-accuracy pseudo-spectral solver (Koehler et al., 2024). These two variants share the same PDE setup and preprocessing but differ in physical parameters: one is more turbulent, whereas the other is comparatively smoother. To emphasize long-horizon robustness, we evaluate on both variants with a closed-loop rollout length of $T=100$ and, for these two datasets, compare DRIFT-Net against scOT under identical settings.
@@
Table 1 reports the final-time relative $L^{1}$ errors across all seven PDE benchmarks under the matched training protocol described above. On the four POSEIDON Navier–Stokes tasks, DRIFT-Net consistently attains the lowest error among all compared architectures, reducing the error of scOT by between $7\%$ and $54\%$ while using fewer parameters. Among the baseline models, scOT is typically the strongest, whereas FNO underperforms on most unsteady flow tasks.
On all three additional PDE benchmarks (Poisson–Gauss, Allen–Cahn, and Wave–Gauss), DRIFT-Net achieves the best test error, with modest gains on Poisson–Gauss and more pronounced improvements on Allen–Cahn and Wave–Gauss. These results indicate that replacing windowed-attention blocks with DRIFT blocks improves accuracy not only on Navier–Stokes benchmarks but also on elliptic, parabolic, and hyperbolic PDEs.
Table 1: Final-time relative $L^{1}$ error (lower is better). Closed-loop rollouts to the common target time $T^{*}$ on all benchmarks. Values are test-set means. Best in bold.
Task scOT FNO U-NO U-Net–ConvNeXt Refiner U-Net DRIFT-Net
NS-SL 3.96 3.69 3.57 3.89 3.48 3.40
NS-PwC 2.35 4.57 1.62 2.22 1.95 1.09
NS-Tracer-PwC 5.18 9.46 10.15 9.63 4.59 4.19
FNS-KF 4.65 4.43 11.70 5.94 4.38 4.32
Poisson–Gauss 0.87 1.29 1.76 1.92 0.86
Allen–Cahn 1.10 1.53 1.27 1.61 1.11 1.03
Wave–Gauss 12.34 16.97 13.94 16.02 13.33 12.15
Long-horizon behavior.
Although final-time metrics are informative, a full closed-loop trajectory over a long horizon provides additional insight into cumulative drift. We therefore analyse two ApeBench-generated forced Kolmogorov variants constructed with a high-accuracy pseudo-spectral solver (Koehler et al., 2024). Both datasets share the same PDE setup but differ in physical parameters, with one being more turbulent and the other comparatively smoother. For both variants we run closed-loop rollouts up to T=100T{=}100 and compare DRIFT-Net against scOT under identical settings.
Table 2 summarizes the final-time relative $L^{1}$ error at T=100T{=}100. On the turbulent variant, DRIFT-Net attains a lower final error than scOT (110.87 vs. 114.14). On the smoother variant, the improvement is more pronounced (57.17 vs. 62.99), indicating increased robustness across different flow regimes.
Table 2: Forced Kolmogorov flows from ApeBench. Final-time relative $L^{1}$ error (lower is better) at T=100T{=}100. Values are test-set means. Best in bold.
Task scOT DRIFT-Net
@@
Beyond the final-time value, we also consider the mean relative $L^{1}$ error across the rollout and a linear growth slope obtained by a least-squares fit of error versus time. On the turbulent variant, DRIFT-Net attains a mean error of $69.89$ compared with $74.76$ for scOT, and an error-growth slope of $1.154$ compared with $1.185$ for scOT. On the smoother variant, the mean error is $31.07$ compared with $35.41$ for scOT, and the slope is $0.581$ compared with $0.641$ for scOT. Figure 2 shows the corresponding error–time curves, which remain consistently lower for DRIFT-Net with slightly flatter tails at large $t$.
Finally, to better visualise how errors accumulate, we plot in Figure 3 the instantaneous error growth on the smoother Kolmogorov variant, defined as the difference between consecutive relative $L^{1}$ values. DRIFT-Net exhibits smaller instantaneous growth than scOT for most of the rollout, especially in the early and mid-range time steps, and only converges to similar values near the end of the horizon. This pattern is consistent with the design goal of DRIFT blocks, namely strengthening global low-frequency coupling while maintaining high-frequency fidelity and avoiding rapid error amplification.
@@
Figure 3: Instantaneous error growth on the smoother ApeBench Kolmogorov dataset. We plot the per-step change in relative $L^{1}$ error over a rollout of length T=100T{=}100. Solid line: DRIFT-Net; dashed line: scOT.
@@
[…]
Table 3: Computational cost on FNS-KF. Parameter count, peak training memory, training time per epoch, and inference throughput (higher is better) for scOT and DRIFT-Net at two width settings. Measurements are taken under identical inference and training settings.
Model Params (M) Peak train mem (GB) Train time / epoch (min) Throughput (steps/s)
scOT 20 17.25 17.30 118
[…]
DRIFT-Net 17 10.87 14.91 158
[…]
@@
We assess the contribution of each major component of DRIFT-Net: Low-Frequency Mixing (LFM), Radial Gating (RG), Frequency-Weighted Loss (FWL), and the two operators in the image branch, namely the $3{\times}3$ depth-wise convolution and the $1{\times}1$ pointwise linear layer. We train ablated variants on NS-PwC and NS-Tracer-PwC, holding all other hyperparameters fixed and using the same training budget and evaluation protocol as the full model. Removing LFM or RG noticeably worsens final-time accuracy; removing FWL also degrades performance, although to a smaller degree. Dropping the $3{\times}3$ depth-wise convolution or the $1{\times}1$ pointwise linear layer further increases the error, showing that the ConvNeXt-style image branch is also important (Table 4). Concretely, relative to the full model, removing LFM increases the final error by $0.56$ and $2.13$ on NS-PwC and NS-Tracer-PwC, respectively; removing RG increases it by $0.61$ and $2.36$; removing FWL yields smaller increases of $0.27$ and $1.17$; removing the depth-wise convolution adds $0.02$ and $0.17$; and removing the pointwise linear layer adds $0.22$ and $0.02$.
@@
w/o $3{\times}3$ depth-wise conv 1.11 4.36
w/o $1{\times}1$ pointwise linear 1.31 4.21
@@ 6 Limitations and Future Work
DRIFT-Net has limitations. It uses a mask to select low frequencies. The initial band is hand-tuned. The mask learns weights later, but the cutoff can be task specific. The model applies FFTs in many blocks. This adds memory traffic and latency on very large grids. Our tests target two-dimensional flow. Three-dimensional flow and coupled multi-physics may bring new training issues and higher cost. Future work will learn spectral partitions end to end. We will study 3D and multi-physics PDEs. We will pair the model with adaptive resolution and mesh refinement. We will also test irregular domains and complex boundary conditions.
@@ 7 Conclusion
We introduced DRIFT-Net, a spectral–spatial coupled integral-kernel neural-operator backbone. Each DRIFT block combines a spectral branch with controlled low-frequency mixing and a ConvNeXt-style image branch, and fuses them via bandwise radial gating in a non-expansive way. A frequency-weighted loss further mitigates spectral bias. On four POSEIDON Navier–Stokes benchmarks, DRIFT-Net reduces final-time relative $L^{1}$ error by $7\%\,$$54\%$ compared to scOT, while using about $15\%$ fewer parameters and achieving higher inference throughput. Additional results on Poisson–Gauss, Allen–Cahn, and Wave–Gauss show that these gains extend beyond Navier–Stokes to elliptic, parabolic, and hyperbolic PDEs. Because DRIFT blocks are modular and architecture-agnostic, they can replace windowed attention blocks in existing multi-scale operator backbones and may be integrated into future PDE foundation models.
Acknowledgments
This project is supported by the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005).
This research includes computations using the computational cluster Katana supported by Research Technology Services at UNSW Sydney.
We also thank Hira Saleem for helpful discussions, feedback, and support during this project; her contributions are gratefully acknowledged.